# SPRec: Leveraging Self-Play to Debias Preference Alignment for Large Language Model-based Recommendations

### v1

@@ Abstract.
Large language models (LLMs) have attracted significant attention in recommendation systems. Current LLM-based recommender systems primarily rely on supervised fine-tuning (SFT) to train the model for recommendation tasks. However, relying solely on positive samples limits the model’s ability to align with user satisfaction and expectations. To address this, researchers have introduced Direct Preference Optimization (DPO), which explicitly aligns recommendations with user preferences using offline preference ranking data. Despite its advantages, our theoretical analysis reveals that DPO inherently biases the model towards a few items, exacerbating the filter bubble issue and ultimately degrading user experience. In this paper, we propose SPRec, a novel self-play recommendation framework designed to mitigate over-recommendation and improve fairness without requiring additional data or manual intervention. In each self-play iteration, the model undergoes an SFT step followed by a DPO step, treating offline interaction data as positive samples and the predicted outputs from the previous iteration as negative samples. This effectively re-weights the DPO loss function using the model’s logits, adaptively suppressing biased items. Extensive experiments on multiple real-world datasets demonstrate SPRec’s effectiveness in enhancing recommendation accuracy and addressing fairness concerns.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Recently, large language models (LLMs) have demonstrated significant success across numerous domains, showcasing advanced capabilities in learning, reasoning, and generalizing to downstream tasks (Zhao et al., 2023; Wang et al., 2024a; Dai et al., 2024). In the field of recommender systems, there has been growing interest in leveraging the potential of LLMs in novel ways (Wu et al., 2024c). One prominent approach involves positioning LLMs as the central recommendation backbone, utilizing users’ past interactions and current needs to generate personalized item recommendations (Bao et al., 2023a; Wei et al., 2024). Compared to traditional methods, LLM-based recommendation systems (LRSs) offer distinct advantages, including a deeper contextual understanding and the flexibility to dynamically adapt to users’ evolving preferences.
@@
To enable LLMs to learn collaborative filtering signals and effectively perform specific recommendation tasks, a prevalent strategy is to fine-tune pre-trained LLMs via Supervised Fine-Tuning (SFT) on offline recommendation datasets (Bao et al., 2023a, b; Liao et al., 2024b; Zhang et al., 2023b). This approach allows LLMs to efficiently internalize user preferences within the offline data by adjusting their parameters to align with the recommendation task. Building on the foundation of SFT, recent research has adopted Direct Preference Optimization (DPO) (Rafailov et al., 2024) to further refine user preferences (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024a). While SFT relies solely on desirable answers, DPO incorporates both chosen and rejected response pairs, allowing the LLM to learn user ranking preferences and develop a more nuanced understanding of fine-grained, personalized information. This approach aligns with the common practice in recommendation models, which utilize both positive and negative samples to train the model effectively (Chen et al., 2023b; Shi et al., 2023).
@@
Despite these advancements, we identify that employing DPO to align user preferences in recommender systems inherently introduces significant biases due to its underlying mechanisms. These biases can lead to serious homogeneity issues, where LLMs produce items with similar names or content. Fig. 1 provides a simplified visualization of how token-level and item-level biases manifest in the Top-$K$ movie recommendation results. Token-level biases occur due to the fact that LLMs generate item names in a tokenized fashion. Since LLMs are usually tuned to maximize the likelihood of target tokens, it can result in an overemphasis on items composed of more common tokens. For instance, movies with common words like “the” in their titles may be disproportionately recommended, regardless of their actual relevance to user preferences. At the item level, biases can emerge from multiple factors, particularly after fine-tuning, where LLMs may disproportionately recommend popular items, such as the Batman film series. This phenomenon can lead to the creation of filter bubbles (Gao et al., 2023a; Wang et al., 2022), where users are repeatedly exposed to a narrow range of highly popular content, ultimately detracting from the diversity of their recommendations and the overall user experience.
@@
Some research has been conducted on bias and unfairness issues in LRSs (Dai et al., 2024; Zhang et al., 2023a; Wang et al., 2024b; Gallegos et al., 2024; Sakib and Das, 2024a). Dai et al. (2024) provided a comprehensive overview of the various types of biases that emerge across different stages of these models and outlined strategies to mitigate them. For example, Jiang et al. (2024) proposed re-weighting the fine-tuning loss for each item and re-ranking the generated results to ensure equal treatment of genre groups. Similarly, Bao et al. (2024) adjusted the LLM decoding process by removing the length normalization term and incorporating predictions from a text-free model, helping to reduce amplification bias and address homogeneity issues. However, these existing methods either require carefully crafted rules or rely heavily on external knowledge, limiting their applicability in general recommendation systems.
@@
To this end, we propose a self-play recommendation tuning framework, SPRec, to adaptively suppress biases and improve fairness in LRSs without the need for additional information or expert knowledge. The core idea of SPRec is very straightforward: each tuning iteration begins with an SFT round using positive samples from offline data, followed by a DPO round still treating SFT data as positive samples and the predicted outputs from the previous iteration as negative samples. The philosophy is to let the model “play” with its own output by re-weighting the DPO loss function based on the predictions. Hence, the higher an item ranks in the model’s predictions, the more it is penalized. Meanwhile, the SFT process reinforces the ranking of positive items. Over time, this self-play learning process adaptively suppresses all undesirable items (biases) while maintaining alignment with positive samples. Extensive experiments on public datasets demonstrate that SPRec effectively enhances both accuracy and fairness metrics, showcasing its potential as a practical and efficient solution for LRSs.
@@
  • We provide an analysis that shows current LRSs tuned through DPO inevitably exhibit biases due to their underlying learning mechanisms, leading to the homogeneity issue.
  • We propose a self-play recommendation tuning framework, SPRec, which addresses these biases and improves fairness without the need for external knowledge.
  • Experiments validate SPRec’s effectiveness in improving accuracy, diversity, and fairness, with ablation studies attributing the improvements to the self-play negative samples.
@@ 2.1. LLMs for Recommendation
LLMs have demonstrated exceptional generative, generalization, and reasoning capabilities in natural language processing, prompting researchers in the recommendation system field to explore their application for personalized recommendations. There are several paradigms for leveraging LLMs in recommendation tasks: (1) using them directly as decision makers (Bao et al., 2023b, 2024), (2) assisting the recommendation process by providing additional information (Sheng et al., 2024; Liu et al., 2024; Geng et al., 2024), and (3) acting as user simulators (Zhang et al., 2024b; Cai et al., 2024; Zhang et al., 2024a). Early research focused on crafting prompts to stimulate the capabilities of LLMs to handle recommendation tasks (Gao et al., 2023b; Hou et al., 2024; Wang et al., 2023b; Gao et al., 2024b).
Subsequently, fine-tuning methods emerged, showing that adapting LLM parameters on recommendation data significantly improves their performance on these tasks. These approaches often rely on Supervised Fine-Tuning (SFT) (Liu et al., 2023; Chen et al., 2025). To further align LLMs with personalized human preferences, researchers adopted Direct Preference Optimization (DPO) for post-training (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024a). However, no prior work has highlighted the inherent mechanisms of DPO that can introduce severe biases in recommender systems. These biases result in recommendations favoring only the exposed items, ultimately harming user experience. In this work, we are the first to identify this issue and propose a mitigation strategy.
@@ 2.2. Biases in Recommender Systems
Bias and unfairness issues are pervasive in recommender systems and have been extensively studied by researchers. Chen et al. (2023a) provide a comprehensive survey summarizing various types of biases, including popularity bias, selection bias, and position bias, among others. When practical recommender systems suffer from these biases, their output can significantly impact user satisfaction. For instance, they may promote content with clickbait titles (Wang et al., 2021) or confine users within filter bubbles, leading to boredom and reduced engagement (Anwar et al., 2024; Areeb et al., 2023; Gao et al., 2023c). Moreover, algorithmic decisions made by recommender platforms can be biased towards certain items, raising concerns about fairness (Wang et al., 2023c; Li et al., 2023). Such biases can disproportionately affect specific user groups or diminish creators’ motivation to produce content (Yao et al., 2024; Jagadeesan et al., 2024).
These issues have persisted throughout the evolution of recommendation systems and continue to exist in the era of LLMs (Wu et al., 2024b; Xu et al., 2024; Tommasel, 2024). Many researchers have identified that LLMs can exhibit social biases, leading to unfair recommendations associated with sensitive attributes such as gender and race (Zhang et al., 2023a; Sakib and Das, 2024b; Shen et al., 2023; Deldjoo and Di Noia, 2024; Ferrara, 2023). Dai et al. (2024) provided a comprehensive review of bias and fairness issues in information retrieval, proposing a unified distribution mismatch perspective to address all forms of bias and unfairness. They also summarized efforts to mitigate bias in LRSs. However, existing efforts either rely on a predefined target distribution or require additional information to guide the alignment direction for LLMs. In this work, we are the first to explore a self-play framework for mitigating bias in LRSs. Unlike previous methods, our approach requires neither prior knowledge nor additional models. By simply altering the tuning procedure, we effectively reduce the long-tail effect and enhance both diversity and fairness in LRSs.
@@ 2.3. Self-Play Mechanism
Machine learning models are often data-driven, relying heavily on the availability of offline data. However, offline data is inherently limited, raising an important question: can algorithms improve themselves iteratively without the need for additional data? This is precisely the challenge that the self-play mechanism aims to address. The concept of self-play originated from board games such as Go and chess (Samuel, 1959; Tesauro et al., 1995), exemplified by groundbreaking systems like AlphaGo Zero (Silver et al., 2017b) and AlphaZero (Silver et al., 2017a).
In the era of LLMs, early preference alignment algorithms like RLHF and DPO operate as single optimization procedures. Building on this, Chen et al. (2024a) proposed that LLMs can refine their capabilities through self-play by interacting with instances of themselves. More specifically, the LLM generates its own training data from prior iterations and subsequently learns a new policy that outperforms the old one (Wu et al., 2024a). This iterative fine-tuning framework has been shown to be equivalent to finding the Nash equilibrium of a two-player game (Munos et al., 2023; Rosset et al., 2024; Calandriello et al., 2024), gaining significant recognition due to its solid theoretical foundation and simplicity (Gao et al., 2024a; Ye et al., 2024). In recommender systems, we leverage the self-play mechanism to adaptively reduce biased items in a simple and effective manner.
@@ 3. Preliminary
Figure 2. Illustration of the pipelines and principles of SFT, DPO, and our proposed SPRec in LLM-based recommendations.
@@ 3.1. Supervised Fine-tuning (SFT)
To enable an open-sourced LLM to effectively learn recommendation tasks, one of the most practical approaches is fine-tuning all or a subset of its parameters using a set of demonstration data derived from offline recommendation logs. The fine-tuning process aligns the model’s behavior with the recommendation objective by optimizing its parameters on the given dataset. The objective function for this fine-tuning process is defined as maximizing the log-likelihood over the offline training dataset $\mathcal{D}$:
@@
where $(x_{i},y_{i})$ denotes the input-output pairs sampled from the training dataset $\mathcal{D}$. Here, $x_{i}$ represents the context of a recommendation, typically comprising user features and a sequence of historically interacted items, while $y_{i}$ denotes the target item to be recommended. Let $p_{\mathcal{D}}(y|x)$ be the probability of $y$ given $x$ in the dataset $\mathcal{D}$, i.e., the popularity. SFT’s formulation is equal to learning to align the model’s predictions with it by minimizing the forward KL-divergence $\mathbb{D}_{KL}(p_{\mathcal{D}}(y|x),\pi_{\theta}(y|x))$:
@@
(2) $=\arg\min_{\pi_{\theta}}\mathbb{E}_{(x_{i},y_{i})\sim D}-\log\pi_{\theta}(y_{i}|x_{i})+H(p_{\mathcal{D}}),$
where $H(p_{\mathcal{D}})$ denotes the entropy of $p_{\mathcal{D}}$, which is constant with respect to the optimization of the LLM parameters $\theta$.
However, relying exclusively on subsequent targets as positive samples may confuse the model about the true objectives. In recommendation tasks, negative samples play a critical role in highlighting undesirable outcomes (Shi et al., 2023). Thus, it is natural to consider incorporating both positive and negative samples from the offline data to better align the model with user preferences.
3.2. RLHF and DPO
To ensure that outputs align with intricate user preferences (such as helpfulness and harmlessness (Bai et al., 2022)), researchers have introduced Reinforcement Learning with Human Feedback (RLHF) to train LLMs using reward signals (Ouyang et al., 2022). Typically, the RLHF stage follows the SFT stage, further refining language models to better adhere to human instructions. RLHF involves learning a reward model $r_{\phi}(x,y)$, which captures human preferences. The optimization objective during this stage is defined as:
(3) $\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}[r_{\phi}(x,y)]-\beta\mathbb{D}_{KL}[\pi_{\theta}(y|x),\pi_{\mathrm{ref}}(y|x)],$
where $\beta$ is a regularization term that penalizes deviations of the learned model $\pi_{\theta}$ from the reference policy $\pi_{\mathrm{ref}}$.
While RLHF has demonstrated remarkable success in aligning LLMs with human preferences, it is computationally intensive, complex, and heavily reliant on the reward model for effectiveness. To simplify reward modeling and mitigate the instability associated with RLHF training, researchers have proposed Direct Preference Optimization (DPO) (Rafailov et al., 2024), which utilizes a closed-form solution of RLHF’s optimal policy. The resulting objective function for DPO is:
@@ 3.2. RLHF and DPO
where $(x,y_{w},y_{l})$ represents a prompt $x$ alongside a chosen (preferred) answer $y_{w}$ and a rejected answer $y_{l}$. In the context of recommendation tasks, $x$ corresponds to the recommendation context, as in the SFT stage, while $y_{w}$ and $y_{l}$ denote a positive and a negative sample, respectively.
DPO offers an efficient and stable alternative to RLHF, aligning models with user preferences while reducing reliance on complex reward modeling. Its ability to naturally leverage positive and negative samples makes it particularly well-suited for recommender systems (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024a).
3.3. Evaluating Biases and Unfairness through Distribution Alignment Perspective

### v3

# SPRec: Self-Play to Debias LLM-based Recommendation

@@ Abstract.
Large language models (LLMs) have attracted significant attention in recommendation systems. Current work primarily applies supervised fine-tuning (SFT) to adapt the model for recommendation tasks. However, SFT on positive examples only limits the model’s ability to align with user preference. To address this, researchers recently introduced Direct Preference Optimization (DPO), which explicitly aligns LLMs with user preferences using offline preference ranking data. However, we found that DPO inherently biases the model towards a few items, exacerbating the filter bubble issue and ultimately degrading user experience.
In this paper, we propose SPRec, a novel self-play framework designed to mitigate over-recommendation and improve fairness without requiring additional data or manual intervention. In each self-play iteration, the model undergoes an SFT step followed by a DPO step, treating offline interaction data as positive samples and the predicted outputs from the previous iteration as negative samples. This effectively re-weights the DPO loss function using the model’s logits, adaptively suppressing biased items. Extensive experiments on multiple real-world datasets demonstrate SPRec’s effectiveness in enhancing recommendation accuracy and fairness. The code is available via https://github.com/RegionCh/SPRec.
@@
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714524isbn: 979-8-4007-1274-6/25/04ccs: Information systems Recommender systems
@@ 1. Introduction
Recently, large language models (LLMs) have demonstrated significant success across numerous domains, showcasing advanced capabilities in learning, reasoning, and generalizing to downstream tasks (Wang et al., 2024b; Dai et al., 2024). In the field of recommender systems, there has been growing interest in leveraging the potential of LLMs (Wu et al., 2024). One prominent approach involves positioning LLMs as the central recommendation backbone, utilizing users’ past interactions and current needs to generate personalized recommendations (Bao et al., 2025; Wei et al., 2024). Compared to traditional methods, LLM-based recommendation systems (LRSs) offer distinct advantages, including a deeper contextual understanding and the flexibility to adapt to users’ evolving preferences.
@@
To enable LLMs to learn collaborative filtering signals and effectively perform item recommendations, a prevalent strategy is to fine-tune pre-trained LLMs via Supervised Fine-Tuning (SFT) (Bao et al., 2025). This approach allows LLMs to efficiently internalize user preferences from offline data by adjusting their parameters to align with the recommendation task. Building on SFT, recent research has adopted Direct Preference Optimization (DPO) (Rafailov et al., 2024) to further refine user preferences (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024). While SFT relies solely on desirable answers, DPO incorporates both chosen and rejected response pairs, allowing the LLM to learn user ranking preferences and gain a more nuanced understanding of fine-grained, personalized information. This approach mirrors the common practice in recommendation models, which utilize both positive and negative samples for effective training (Chen et al., 2023b; Shi et al., 2023).
@@
Despite these advancements, we find that employing DPO to align user preferences in recommender systems inherently introduces significant biases due to its underlying mechanisms. These biases can lead to serious homogeneity issues, where LLMs recommend items with similar names or content. Fig. 1 illustrates how token-level and item-level biases manifest in the Top-$K$ movie recommendation. Token-level biases arise due to the fact that LLMs generate item names in a tokenized fashion. Since LLMs are usually tuned to maximize the likelihood of target tokens, items with more common tokens (e.g., movies with the word “the” in their titles) may be overrepresented, regardless of user relevance. At the item level, biases can emerge from multiple factors, particularly after fine-tuning, where LLMs may disproportionately recommend popular items, such as the Batman film series. This can lead to filter bubbles (Gao et al., 2023a, c), where users are repeatedly exposed to a narrow range of popular content, limiting the diversity of recommendations and degrading the user experience.
@@
Some research has been conducted on bias and unfairness issues in LRSs (Dai et al., 2024; Zhang et al., 2023; Gallegos et al., 2024). Dai et al. (2024) provided a comprehensive overview of the various types of biases that emerge across different stages of these models and outlined strategies to mitigate them. For example, Jiang et al. (2024) proposed re-weighting the fine-tuning loss for each item and re-ranking the generated results to ensure equitable treatment across genre groups. Similarly, Bao et al. (2024) adjusted the LLM decoding process by removing the length normalization term and incorporating predictions from a text-free model, helping to reduce amplification bias and address homogeneity issues. However, these methods often rely on carefully crafted rules or external knowledge, limiting their broader applicability in general recommendation systems.
@@
To this end, we propose a self-play recommendation tuning framework, SPRec, to adaptively suppress biases and improve fairness in LRSs without the need for additional data or expert knowledge. The core idea of SPRec is straightforward: each tuning iteration begins with an SFT round using positive samples from offline data, followed by a DPO. In the DPO step, the SFT data is treated as positive samples, while the predicted outputs from the previous iteration are treated as negative samples. The philosophy is to let the model “play” with its own output by re-weighting the DPO loss function based on its predictions. As a result, items that rank higher in the model’s predictions are penalized, while the SFT process reinforces the ranking of positive items. Over time, this self-play learning process adaptively suppresses undesirable items (biases) while maintaining alignment with positive samples. Extensive experiments on public datasets demonstrate that SPRec effectively improves both accuracy and fairness, showcasing its potential as a practical and efficient solution for LRSs.
@@
  • We analyze how current LRSs tuned through DPO inevitably exhibit biases due to their underlying learning mechanisms, leading to the homogeneity issue.
  • We propose SPRec, a self-play recommendation tuning framework that addresses these biases and improves fairness without the need for external knowledge.
  • Experiments validate SPRec improves accuracy, diversity, and fairness, with ablation studies indicating that the self-play negative samples contribute significantly to the improvements.
@@ 2.1. LLMs for Recommendation
LLMs have shown exceptional generative, generalization, and reasoning capabilities in NLP, driving research into their applications for personalized recommendations. Their integration into recommendation tasks follows three main paradigms: (1) acting as decision makers (Bao et al., 2024; Mao et al., 2024), (2) assisting by providing contextual information (Liu et al., 2024; Geng et al., 2024), and (3) serving as user simulators (Zhang et al., 2024; Cai et al., 2024). Early studies explored prompt engineering to leverage LLMs for recommendation tasks (Gao et al., 2023b; Hou et al., 2024).
Later, fine-tuning methods emerged, demonstrating that adapting LLM parameters on recommendation data significantly enhances performance. These approaches primarily rely on SFT (Bao et al., 2024; Chen et al., 2025). To further align LLMs with user preferences, DPO has been employed for post-training (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024). However, prior work has overlooked DPO’s inherent tendency to introduce severe biases, favoring only frequently exposed items and degrading user experience. In this work, we are the first to identify this issue and propose a mitigation strategy.
@@ 2.2. Biases in Recommender Systems
Bias and fairness issues are pervasive in recommender systems and have been extensively studied. Chen et al. (2023a) provide a comprehensive survey on biases such as popularity bias, selection bias, and position bias. These biases can significantly impact user satisfaction, promoting clickbait content or reinforcing filter bubbles that reduce engagement (Gao et al., 2023c). Additionally, algorithmic decisions may favor certain items, raising fairness concerns (Wang et al., 2023; Li et al., 2023), disproportionately affecting user groups and discouraging content creators (Yao et al., 2024; Jagadeesan et al., 2024).
These challenges persist in LLM-based recommender systems (Wu et al., 2024; Xu et al., 2024; Tommasel, 2024). Research shows that LLMs can inherit social biases, leading to unfair recommendations related to sensitive attributes like gender and race (Zhang et al., 2023). Dai et al. (2024) provide a unified distribution mismatch perspective on bias and fairness in information retrieval. Existing bias mitigation methods in LRSs typically rely on predefined target distributions or external guidance for LLM alignment. In contrast, we introduce the first self-play framework for mitigating bias in LRSs, requiring neither prior knowledge nor additional models. By simply modifying the tuning process, our approach reduces long-tail effects and improves fairness.
@@ 2.3. Self-Play Mechanism
Machine learning models are often data-driven, relying heavily on the availability of offline data. However, offline data is inherently limited, raising an important question: can algorithms improve themselves iteratively without the need for additional data? This is precisely the challenge that the self-play mechanism aims to address. The concept of self-play originated from board games such as Go and chess, exemplified by groundbreaking systems like AlphaGo Zero (Silver et al., 2017) and AlphaZero (Silver et al., 2018).
In the era of LLMs, early preference alignment algorithms like RLHF and DPO operate as single optimization procedures. Building on this, Chen et al. (2024a) proposed that LLMs can refine their capabilities through self-play by interacting with instances of themselves. More specifically, the LLM generates its own training data from prior iterations and subsequently learns a new policy that outperforms the old one (Wu et al., 2025). This iterative fine-tuning framework has been shown to be equivalent to finding the Nash equilibrium of a two-player game, gaining significant recognition due to its solid theoretical foundation and simplicity (Calandriello et al., 2024). In recommender systems, we leverage the self-play mechanism to adaptively reduce biased items in a simple and effective manner.
@@ 3.1. Supervised Fine-tuning (SFT)
To enable an open-source LLM to learn recommendation tasks effectively, a practical approach is fine-tuning all or part of its parameters using demonstration data from offline recommendation logs. The objective is to align the model’s behavior with the recommendation task by maximizing the log-likelihood over the training dataset $\mathcal{D}$:
@@
where $(x_{i},y_{i})$ are input-output pairs from $\mathcal{D}$, with $x_{i}$ representing user context and interaction history, and $y_{i}$ the target item. Defining $p_{\mathcal{D}}(y|x)$ as the empirical probability (i.e., item popularity), SFT aligns model predictions by minimizing the forward KL-divergence:
@@
(2) $=\arg\min_{\pi_{\theta}}\mathbb{E}_{(x_{i},y_{i})\sim\mathcal{D}}-\log\pi_{\theta}(y_{i}|x_{i})+H(p_{\mathcal{D}}),$
where $H(p_{\mathcal{D}})$ is the constant entropy of $p_{\mathcal{D}}$.
3.2. Direct Preference Optimization (DPO)
To ensure that model outputs align with intricate user preferences, researchers have proposed Direct Preference Optimization (DPO) (Rafailov et al., 2024), which optimizes the following objective function:
@@ 3.2. Direct Preference Optimization (DPO)
where $(x,y_{w},y_{l})$ denotes a prompt $x$ with a chosen (preferred) answer $y_{w}$ and a rejected answer $y_{l}$. The parameter $\beta$ acts as a regularization factor, controlling the extent to which the learned policy $\pi_{\theta}$ deviates from the reference policy $\pi_{\mathrm{ref}}$.
In the context of recommendation tasks, $x$ represents the user context, typically comprising user features and historical interaction sequences, while $y_{w}$ and $y_{l}$ correspond to positive and negative samples, respectively. The goal is to encourage the model to assign higher probabilities to preferred items ($y_{w}$) over less desirable ones ($y_{l}$), effectively capturing user preferences.
DPO offers an efficient and stable solution for preference alignment, eliminating the need for complex reward models often required in reinforcement learning-based approaches. Its natural ability to incorporate both positive and negative samples makes it particularly well-suited for recommender systems, where learning from contrasting user interactions is crucial (Chen et al., 2024b; Bai et al., 2024; Liao et al., 2024).
3.3. Evaluating Bias via Distribution Alignment